OpenStudy (anonymous):

The first order derivative of a function y (x ) is given as y'(x)=3x^4-5x^2 If the function y(x) passes through the origin i.e. (0,0), determine y(x) by finding the anti-derivative of y'(x) . Help please?

OpenStudy (anonymous):

Ok, so this is some pretty simple integration. Do you know how to do integration by parts?

OpenStudy (anonymous):

no because this new chapter...so I ned help to solve it

OpenStudy (anonymous):

Alright, no problem. you use the basic formula for integration which is \[\frac{ x^{n-1} }{ n-1 }\]

OpenStudy (cwrw238):

integrating y' gives 3 x5/5 - 5 x^3/3 + C where C is is some constant y = 3 x5/5 - 5 x^3/3 + C plug in y = 0 and x = 0 to find the value of C

OpenStudy (anonymous):

So for that first part of the problem 3x^4 you would subtract 1 from the exponent (the 4) and then divide by that.

OpenStudy (cwrw238):

@stardogchampion - thats differentiation

OpenStudy (anonymous):

Oh my I'm an idiot.

OpenStudy (cwrw238):

Integration is x^(n+1) ------- n + 1

OpenStudy (anonymous):

Haha, sorry.

OpenStudy (cwrw238):

lol - we all are on times....

OpenStudy (cwrw238):

if we solve 3 (0)^5 / 5 - 5 (0)^3 / 3 + C = 0 for C what do we get/

OpenStudy (cwrw238):

C = 0 right?

OpenStudy (anonymous):

yup

OpenStudy (cwrw238):

finally we can write y(x) which is